# Just Ramp-up: Unleash the Potential of Regression-based Estimator for A/B Tests under Network Interference

### v2

@@ 1 Introduction
There has been a series of works studying spatiotemporal interference with panel data [1, 53, 61], which consider estimands different from static setting and usually require stronger assumptions, e.g. factor model. On one hand, there exists a variety of temporal interference in different scenarios [28, 23, 24, 73, 12], while we target developing general methodology. On the other hand, we believe the estimation of static GATE has not yet been solved elegantly, particularly regarding simulation performance. Therefore, we do not consider any temporal interference in the ramp-up process; instead, we focus on network interference and leverage multiple experiments for estimating the static GATE. Nonetheless, our analysis incorporates temporal correlation of treatment assignments, e.g. the staggered rollout design [41, 3, 73], which ensures temporal monotonicity in the treatment of any specific unit and introduces strong temporal correlation in its treatments over different time points. In addition, we mention that our framework can easily accommodate dynamic graph structure and temporal interference, though the latter is concerned with the change of estimand and out of the scope of this paper.
@@
For estimators, beyond classic Horvitz-Thompson (HT) and difference-in-means estimators, there have been lots of estimators targeting GATE estimation under network interference, such as HT-style estimator based on different exposure mappings [64, 22], Hájek estimator [27, 65], cluster-adaptive estimator [49]. In this paper, we focus on the regression-based estimator for static GATE. Besides, [27] discusses the advantages of regression-based estimators, such as convenience for implementation, and capacity to incorporate covariates. Nonetheless, we emphasize another point: regression on network experiment data actually models the interference effect. Our core methodology, merging experiment data at different steps, can significantly improve the training of such regression models since it augments the training data intrinsically. In contrasts, we will demonstrate that repeating the experiment with the same treatment proportion is much less effective for training regression model.
@@
The key methodology proposed and discussed in this paper is merging data from multiple experiments that are selected from the ramp-up process for training a regression model for predicting unit outcomes. We first develop theoretical results on bias and variance of simple linear regression under general linear interference. We demonstrate that bias dominates when the interference is of regular intensity, and our merging procedure can substantially reduce the bias. However, the bias reduction achieved with the linear regression model remains unsatisfactory due to its simplicity. Considering the predominant role of bias, we then explore an enhanced estimator based on GNN. Through extensive simulation studies, we verify that most of the conclusions derived in the context of linear regression still hold for the GNN-based estimator. Specifically, we verify that the GNN estimator trained on the merged data composed of two steps achieves the state-of-the-art mean squared error (MSE) with balanced bias and standard deviation in linear interference setting, as illustrated in theoretical analysis of linear regression estimator. Moreover, we also validate that merging more than two steps of experimental data can be beneficial under nonlinear interference.
@@ 2.2 General Linear Interference Model
$Y(\mathbf{z})=\beta_{0}+\beta_{1}\mathbf{z}+\beta_{2}B\mathbf{z}+\mathbf{\epsilon}$ (2)
Where $\beta_{0},\beta_{1},\beta_{2}$ and $B$ are all unknown parameters, and we assume $\beta_{2}\neq 0$, indicating that there is indeed an interference effect. This model can be written more compactly as:
$Y(\mathbf{z})=\begin{pmatrix}\mathbf{1}&\mathbf{z}&B\mathbf{z}\end{pmatrix}\beta+\mathbf{\epsilon}$ (3)
We further assume simple Gaussian noise structure to facilitate theoretical analysis.
@@
The matrix $B$ in (2) characterizes the interference effects. In this paper, we restrict $B_{ii}=0$ for all $i\in[n]$ to ensure that $\beta_{1}$ and $\beta_{2}$ accurately represent the direct treatment effect and the interference effect, respectively. For simplicity, we do not assume any heterogeneity in the base effect $\beta_{0}$ and the direct effect $\beta_{1}$ among units. Instead, we put our attention to a general interference matrix $B$, which is capable of modeling complex long-distance dependencies in the graph. We provide several specific matrices $B$ in the following, which will serve as running examples throughout this paper.
@@
If we take $B=D^{-1}A$, then our potential outcome model is reduced to the classic linear-in-means model [2], formatted as follows:
$Y_{i}(\mathbf{z})=\beta_{0}+\beta_{1}z_{i}+\beta_{2}\frac{\sum_{j\in\mathcal{N}(i)}z_{j}}{\deg_{i}}+\mathbf{\epsilon}$ (5)
@@
We mention that removing the degree normalization, i.e. setting $B=A$ results in another common case where the exposure condition shifts from the proportion to the number of treated neighbors, thereby characterizing a different class of social interaction.
@@ 3 Analysis on Linear Regression Estimator
In this section, we perform a detailed analysis of the simple linear regression model [^1] We refer to it as ”simple” because it does not incorporate knowledge of graph topology; that is, $f(\mathbf{z},A)=f(\mathbf{z})$ for $f\in\mathcal{F}_{lr}$. As will be demonstrated in the subsequent subsections, this regression model supports precise analysis on bias and variance and provides a clear illustration of the role of merging experimental data for training. Many of the phenomena observed with this simple regression model can also be generalized to other regression models, such as those in $\mathcal{F}_{gnn}$, which will be discussed in the simulation study. On the other hand, we also note that precise statistical analysis would become infeasible if a component of network exposure is involved in the regression model. In summary, the linear regression model examined here is both straightforward and sufficiently powerful for illustration, allowing for the development of insightful conclusions. To be specific, the considered linear regression model is given by:
$Y=X(\mathbf{z})\theta+\mathbf{\epsilon_{\theta}}+\epsilon^{\prime}$ (9)
here $X(\mathbf{z})=(\mathbf{1},\mathbf{z})$, $\theta=(\theta_{0},\theta_{1})^{\top}$, and $\epsilon^{\prime}$ is the error term.
After estimating $\hat{\theta}$ from the data, our GATE estimator is given by:
Linear regression admits a closed-form optimal solution that brings huge technical convenience. Nonetheless, there exists additional complexity lying in the randomness of the design matrix $X(\mathbf{z})$, since it incorporates the randomness of treatment assignment. As is well known, the closed-form solution of ordinary least squares (OLS) is
@@
here we keep the notation $X(\mathbf{z})$ to stress its dependency on experimental design.
The inverse of such a matrix involves a determinant in the denominator, making the analysis intractable if this determinant also incorporates randomness. Therefore, we consider $\dim(\theta)=2$ and complete randomization scheme in the following analysis.
@@ 3.1 One-step Experiment
In this part, we first examine the simplest case in that only one experiment is conducted, i.e. $T=1$. We specify the number of treated units in the experimental data as $d=cn$ and construct the following conclusion:
@@
$\operatorname{Bias}(\hat{\tau})$ $=\mathbb{E}[\hat{\theta}_{1}]-(\beta_{1}+\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B))$ (12)
$=\frac{\beta_{2}}{nc(1-c)}\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-c\mathbf{1})^{\top}\right]B\right)-\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)$
$=-\frac{\beta_{2}}{n}(\sum_{i,j}B_{ij})\left(\frac{1}{n(n-1)}+1\right)$
@@
$\operatorname{Var}(\hat{\tau})/\beta_{2}^{2}$ $=(\frac{1}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)+\frac{1}{\left(\beta_{2}nc(1-c)\right)^{2}}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}(\mathbf{z}-c\mathbf{1})]\sigma^{2}_{e}$ (13)
$=(\frac{1}{nc(1-c)})^{2}\left((\sum_{i,j}B_{ij})^{2}\frac{c(d-1)}{(n-1)}\left(\frac{c(d-n)+6+6c^{2}-12c}{(n-2)(n-3)}\right)\right.$
$+(\sum_{i,j,k}B_{ij}B_{kj})\frac{c(d-n)+2c^{2}+2-4c}{(n-1)(n-2)}$
$+(\sum_{i,j,l}B_{ij}B_{il})\left((1-2c)c\frac{(d-1)(d-2)}{(n-1)(n-2)}+c^{3}\frac{d-1}{n-1}\right)$
$+(\sum_{i,j}B_{ij}^{2})\left((1-2c)c\frac{d-1}{n-1}+c^{3}\right)$
$+\left.(\sum_{i,j}B_{ij}B_{ji})(1-c)^{2}c\frac{d-1}{n-1}\right)$
$-(\sum_{i,j}B_{ij})^{2}\frac{1}{\left(n(n-1)\right)^{2}}$
$+\frac{\sigma_{e}^{2}}{\beta_{2}^{2}nc(1-c)}$
The notation such as $\sum_{i,j}$ denotes summation over $i\in[n],j\in[n]$ without additional constraints. The details of the proof are provided in Appendix A.1. We now proceed to elaborate on the results presented above.
First, we notice that the intensity of direct effect, $\beta_{1}$, does not influence either bias or variance. In fact, only the intensity of interference, $\beta_{2}$, appears in the expressions above. Then we check the expression of bias, where the part from estimator $\mathbb{E}[\hat{\theta}_{1}]$ contributes to the increase of bias, even though its order of quantity is much lower than that of relative bias [^2] Hereafter, we use the term relative bias to denote $\text{Bias}/\beta_{2}$. induced by interference effect. This implies that the primary source of bias is the interference, while interference term is absent in this linear regression.
@@
$\operatorname{Bias}(\hat{\tau})/\beta_{2}$ $=\Theta(1)$ (20)
$\operatorname{Var}(\hat{\tau})/\beta_{2}^{2}$ $=\Theta(1/n)$
@@
We then further clarify the implications of Assumption 1. For a specific unit $i$, the interference on its outcome is represented by $\beta_{2}\sum_{j}B_{ij}z_{j}$, and the treatment effect is $\beta_{1}$. The first part of Assumption 1 restricts the interference effect received by unit $i$ to be of constant order, ensuring that interference does not dominate the outcome for unit $i$. Additionally, this assumption allows for some units to be minimally affected by interference. At last, through basic algebra, one can verify that $\sum_{j}|B_{ij}|$ is equal to $1$ in Example 1 and less than [^3] If we do not remove the diagonal elements of $(D^{-1}A)^{m}$, then we would exactly have $\sum_{j}|B_{ij}|=\sum_{m=1}^{M}r_{m}$. $\sum_{m=1}^{M}r_{m}$ in Example 2.
@@ 3.2 Two-step Experiment
Then we derive following theorem on the bias term.
@@
The bias of linear regression estimator trained on merged data $(T=2)$ under general linear interference is given by
$\operatorname{Bias}(\hat{\tau})/\beta_{2}$ $=-\frac{\sum_{ij}B_{ij}}{n}\left(1-\frac{(n-1)(c_{1}-c_{2})^{2}+2(c_{1}^{2}+c_{2}^{2})-2(c_{1}+c_{2})}{(n-1)(c_{1}+c_{2})(2-(c_{1}+c_{2}))}\right)$ (23)
We maintain the assumption $c_{1},c_{2}\in[0,1]$ and $c_{1}<c_{2}$, as introduced in the Section 2. Through basic algebra, we can verify that
$01-\frac{(n-1)(c_{1}-c_{2})^{2}+2(c_{1}^{2}+c_{2}^{2})-2(c_{1}+c_{2})}{(n-1)(c_{1}+c_{2})(2-(c_{1}+c_{2}))}>$ 0 (24)
which means training on merged experimental data with treatment proportion $c_{1}$ and $c_{2}$ can substantially reduce the bias, and the reduction on relative bias compared to training on single experimental dataset (with treatment proportion $c$) is
$\frac{\sum_{ij}B_{ij}}{n}\left(\frac{(n-1)(c_{1}-c_{2})^{2}+2(c_{1}^{2}+c_{2}^{2})-2(c_{1}+c_{2})}{(n-1)(c_{1}+c_{2})(2-(c_{1}+c_{2}))}+\frac{c}{n(n-1)}\right)$ (25)
If we ignore the $O(1/n)$ terms in the expression above, the substantial bias reduction is
$\frac{\sum_{ij}B_{ij}}{n}\left(\frac{(c_{1}-c_{2})^{2}}{(c_{1}+c_{2})(2-(c_{1}+c_{2}))}\right)$ (26)
@@ 3.3 $T$-step Experiment
For linear regression estimator trained on merged $T$-step data, the bias is given by
$\operatorname{Bias}(\hat{\tau})/\beta_{2}$ $=-\frac{\sum_{ij}B_{ij}}{n}\left(1-\frac{T\sum_{t=1}^{T}c_{t}^{2}-(\sum_{t=1}^{T}c_{t})^{2}}{(\sum_{t=1}^{T}c_{t})(T-\sum_{t=1}^{T}c_{t})}\right)+O(\frac{1}{n})$ (28)
By Cauchy-Schwarz inequality and basic algebra, we can verify that
@@
Merging data from more than two experiments with varying treatment proportions can still lead to substantial bias reduction compared to the single experiment, similar to the case where $T=2$.
@@
$\operatorname{Var}(\hat{\tau})/\beta_{2}^{2}=\Theta(\frac{1}{n})$ (35)
@@
$\operatorname{Var}(\hat{\tau})/\beta_{2}^{2}=\Theta(\frac{1}{n})$ (37)
@@ 5.1 Basic Setting
We begin by outlining the setup of our simulation study. Rather than using a simple random graph, we utilize a Facebook social network FB-Stanford3999The network topology data can be found in https://networkrepository.com/socfb-Stanford3.php [58], which has also been employed in semi-synthetic experiments in [65, 18]. The network topology $\mathcal{G}$ consists of $|\mathcal{V}|=11586$ nodes and $|\mathcal{E}|=568309$ edges, where each node represents a person, and the edges indicate friendships or social relationships. Based on this real network, we generate the potential outcomes with specific models in the subsequent sections to validate the effectiveness of our methodology.

### v3

@@ 1 Introduction
In this paper, we focus on the regression-based estimator for static GATE.  [23] discusses the advantages of regression-based estimators, such as convenience for implementation, and capacity to incorporate covariates. Nonetheless, we emphasize another point: regression on network experiment data actually models the interference effect. Our core methodology, merging experiment data at different steps, can significantly improve the training of such regression models since it augments the training data intrinsically. In contrasts, we will demonstrate that repeating the experiment with the same treatment proportion is much less effective for training regression model.
This paper proposes a method that leverages merged data from multiple experiments conducted during the ramp-up phase to train a regression model for counterfactual prediction. We view the task of GATE estimation as a counterfactual prediction problem between two extreme exposure levels: global treatment and global control, which inevitably involves extrapolation. The core innovation of our method lies in the merging procedure itself, which amplifies variation in units’ treatment exposures. This enhanced variation provides critical benefits for training the regression model, enabling it to generalize more effectively across counterfactual scenarios and resulting in a more precise GATE estimate.
@@
We first develop theoretical results on the bias and variance of a simple linear regression estimator that neglects network interference, while the true mechanism admits general interference that can pervade the whole graph. Our analysis shows that bias is the dominant factor when the interference is of regular intensity and that our merging procedure can substantially reduce this bias. However, the bias reduction achieved with the linear regression model remains unsatisfactory due to its simplicity. Given the predominant role of bias, we then explore a refined estimator based on GNN. Through extensive simulation studies, we verify that most of the conclusions derived in the context of linear regression still hold for the GNN-based estimator. Specifically, we demonstrate that the GNN estimator, trained on merged data from two experimental steps, performs excellently in mean squared error (MSE) with a balanced bias and standard deviation in the linear interference setting, as illustrated in the theoretical analysis of the linear regression estimator. Moreover, we validate that merging data from more than two experimental steps can be beneficial in settings with nonlinear interference.
@@ 2.2 General Linear Interference Model
$Y(\mathbf{z})=\beta_{0}+\beta_{1}\mathbf{z}+B\mathbf{z}+\mathbf{\epsilon}$ (2)
where $\beta_{0},\beta_{1}$ and $B$ are all unknown parameters. We further assume simple Gaussian noise structure to facilitate theoretical analysis.
@@
The matrix $B$ in (2) characterizes the interference effects. In this paper, we restrict $B_{ii}=0$ for all $i\in[n]$ to separate the direct treatment effect and the interference effect.
For simplicity, we do not assume any heterogeneity in the base effect $\beta_{0}$ and the direct effect $\beta_{1}$ among units. Instead, we put our attention to a general interference matrix $B$, which is capable of modeling complex long-distance dependencies in the graph. We provide two specific matrices $B$ in the following, which will serve as running examples throughout this paper.
@@
If we take $B=rD^{-1}A$, where $r$ is the scale parameter, then our potential outcome model is reduced to the classic linear-in-means model [1], formatted as follows:
$Y_{i}(\mathbf{z})=\beta_{0}+\beta_{1}z_{i}+r\frac{\sum_{j\in\mathcal{N}(i)}z_{j}}{\deg_{i}}+\mathbf{\epsilon}$ (4)
@@
We then provide an example that incorporates multihop interference, which demonstrates the generality of this potential outcome model.
@@ 3 Analysis on Linear Regression Estimator
In this section, we analyze the bias and variance of the simple linear regression model that only regress under the general linear network interference introduced in the previous section. As will be demonstrated in the subsequent subsections, this regression model provides a clear illustration of the role of merging experimental data for training. Many of the phenomena observed with this simple regression model can also be generalized to other regression models, such as those in $\mathcal{F}_{gnn}$, which will be shown in the simulation study. In summary, the linear regression model examined here is both technically tractable and sufficiently powerful for illustration, allowing for the development of insightful conclusions. To be specific, we consider the linear regression predictor which does not explicitly account for interference.
$f(\mathbf{z},A)=X(\mathbf{z})\hat{\theta}$ (8)
Here, $X(\mathbf{z})=(\mathbf{1},\mathbf{z})$, $\hat{\theta}$ is the ordinary least squares (OLS) estimator
@@
We keep the notation $X(\mathbf{z})$ to stress its dependency on experimental design. Now the GATE estimator is given by $\hat{\tau}=(\hat{\theta}_{1}+\hat{\theta}_{0})-\hat{\theta}_{0}=\hat{\theta}_{1}$.
Linear regression admits a closed-form optimal solution that brings huge technical convenience. Nonetheless, there exists additional complexity lying in the randomness of the design matrix $X(\mathbf{z})$, since it incorporates the randomness of treatment assignment. The matrix $(X(\mathbf{z})^{\top}X(\mathbf{z}))^{-1}$ involves a determinant in the denominator, making the analysis intractable if the determinant itself is random. This explains why we restrict our consideration to such a simple linear regression estimator and the complete randomization.
@@ 3.1 One-step Experiment
Throughout this section, we use $\hat{\tau}_{(T)}$ to denote the OLS GATE estimator based on $T$-step experiment. In this subsection, we examine the case of $T=1$. We specify the number of treated units in the experimental data as $d=cn$ and derive the bias and variance for $\hat{\tau}_{(1)}$.
@@
$\operatorname{Bias}(\hat{\tau}_{(1)})=-\frac{(\sum_{i,j}B_{ij})}{n}\left(\frac{1}{n(n-1)}+1\right)$ (10)
@@
$\operatorname{Var}(\hat{\tau}_{(1)})$ $=(\frac{1}{nc(1-c)})^{2}\left((\sum_{i,j}B_{ij})^{2}(\frac{c^{3}(c-1)}{n}+O(\frac{1}{n^{2}}))\right.+(\sum_{i,j,k}B_{ij}B_{kj})(\frac{c(c-1)}{n}+O(\frac{1}{n^{2}}))$ (11)
$+(\sum_{i,j,l}B_{ij}B_{il})(c^{3}(1-c)+O(\frac{1}{n}))+(\sum_{i,j}B_{ij}^{2})(c^{2}(1-c)+O(\frac{1}{n}))$
$+\left.(\sum_{i,j}B_{ij}B_{ji})((1-c)^{2}c^{2}+O(\frac{1}{n}))\right)-(\sum_{i,j}B_{ij})^{2}\frac{1}{\left(n(n-1)\right)^{2}}+\frac{\sigma_{e}^{2}}{nc(1-c)}$
Notations like $\sum_{i,j}$ denote summation over $i\in[n],j\in[n]$ without additional constraints. The details of the proof are provided in Appendix A.1. We now proceed to elaborate on the results presented above.
First, we notice that the intensity of direct effect, $\beta_{1}$, does not influence either bias or variance. Then we check the expression of bias, where the part from estimator $\mathbb{E}[\hat{\theta}_{1}]$ contributes to the increase of bias, even though its order of quantity is much lower than that of bias induced by interference effect. This implies that the primary source of bias is the interference, while interference term is absent in this linear regression.
@@
$\operatorname{Bias}(\hat{\tau}_{(1)})$ $=\Theta(1)$ (18)
$\operatorname{Var}(\hat{\tau}_{(1)})$ $=\Theta(1/n)$
@@
We then further clarify the implications of Assumption 1. For a specific unit $i$, the interference on its outcome is represented by $\sum_{j}B_{ij}z_{j}$, and the treatment effect is $\beta_{1}$. The first part of Assumption 1 restricts the interference effect received by unit $i$ to be of constant order, ensuring that interference does not dominate the outcome for unit $i$. Additionally, this assumption allows for some units to be minimally affected by interference. At last, through basic algebra, one can verify that $\sum_{j}|B_{ij}|$ is equal to $1$ in Example 1 and less than [^1] If we do not remove the diagonal elements of $(D^{-1}A)^{m}$, then we would exactly have $\sum_{j}|B_{ij}|=\sum_{m=1}^{M}r_{m}$. $\sum_{m=1}^{M}r_{m}$ in Example 2.
@@ 3.2 Two-step Experiment
We then characterize the bias of the 2-step estimator $\hat{\tau}_{(2)}$ in the following theorem. In the next subsection, we will show that the variance for general $\hat{\tau}_{(T)}$ is persistly negligible. We maintain the assumption $c_{1},c_{2}\in[0,1]$ and $c_{1}<c_{2}$, as introduced in the Section 2.
@@
The bias of the linear regression estimator trained on merged data $(T=2)$ under general linear interference is given by:
$\operatorname{Bias}(\hat{\tau}_{(2)})$ $=-\frac{\sum_{ij}B_{ij}}{n}\left(1-\frac{(n-1)(c_{1}-c_{2})^{2}+2(c_{1}^{2}+c_{2}^{2})-2(c_{1}+c_{2})}{(n-1)(c_{1}+c_{2})(2-(c_{1}+c_{2}))}\right)$ (21)
When $n>2(c_{1}+c_{2}-c_{1}^{2}-c_{2}^{2})/(c_{1}-c_{2})^{2}+1$, we have:
$\frac{(n-1)(c_{1}-c_{2})^{2}+2(c_{1}^{2}+c_{2}^{2})-2(c_{1}+c_{2})}{(n-1)(c_{1}+c_{2})(2-(c_{1}+c_{2}))}\in(0,1)$ (22)
This directly implies that $\operatorname{Bias}(\hat{\tau}_{(1)})$ and $\operatorname{Bias}(\hat{\tau}_{(2)})$ shares the same sign. A substantial reduction in the magnitude of bias from $\hat{\tau}_{(1)}$ to $\hat{\tau}_{(2)}$ is given by:
$|\operatorname{Bias}(\hat{\tau}_{(1)})|-|\operatorname{Bias}(\hat{\tau}_{(2)})|=\frac{|\sum_{ij}B_{ij}|}{n}\frac{(c_{1}-c_{2})^{2}}{(c_{1}+c_{2})(2-(c_{1}+c_{2}))}+O(\frac{1}{n})$ (23)
@@ 3.3 $T$-step Experiment
For linear regression estimator trained on merged $T$-step data, the relative bias is given by:
$\operatorname{Bias}(\hat{\tau}_{(T)})$ $=-\frac{\sum_{ij}B_{ij}}{n}\left(1-\frac{T\sum_{t=1}^{T}c_{t}^{2}-(\sum_{t=1}^{T}c_{t})^{2}}{(\sum_{t=1}^{T}c_{t})(T-\sum_{t=1}^{T}c_{t})}\right)+O(\frac{1}{n})$ (25)
Using the same logic as in Theorem 1, by Cauchy-Schwarz inequality, we can verify:
@@
Therefore, merging data from $T$-step experiments with varying treatment proportions can still achieve substantial bias reduction compared to using a single experiment.
@@
$\operatorname{Var}(\hat{\tau}_{(T)})=\Theta(\frac{1}{n})$ (32)
@@
$\operatorname{Var}(\hat{\tau}_{(T)})=\Theta(\frac{1}{n})$ (34)
@@ 5.1 Basic Setting
We begin by outlining the setup of our simulation study. Rather than using a simple random graph, we utilize a Facebook social network FB-Stanford3777The network topology data can be found in https://networkrepository.com/socfb-Stanford3.php [51], which has also been employed in semi-synthetic experiments in [57, 15]. The network topology $\mathcal{G}$ consists of $|\mathcal{V}|=11586$ nodes and $|\mathcal{E}|=568309$ edges, where each node represents a person, and the edges indicate friendships or social relationships. Based on this real network, we generate the potential outcomes with specific models in the subsequent sections to validate the effectiveness of our methodology.